OpenStudy (anonymous):

log8x - log (1+ x^1/2) = 2

OpenStudy (anonymous):

do you know how to reduced the difference between to logs?

OpenStudy (kira_yamato):

Try combining the 2 logarithms together. It helps a lot :)

OpenStudy (kira_yamato):

But keep in mind that x > 0

OpenStudy (unklerhaukus):

is 8 the base of the first log?

OpenStudy (unklerhaukus):

or are both logs to base 10?

OpenStudy (anonymous):

by division? would it be 8x/(1+x^1/2)=100. if its so. what next? What do I do next! yes

OpenStudy (anonymous):

both are base 10

OpenStudy (kira_yamato):

Get rid of the denominator.

OpenStudy (anonymous):

how?

OpenStudy (kira_yamato):

Multiply the equation throughout by the denominator

OpenStudy (anonymous):

I did that! seems to make it more complicated.

OpenStudy (kira_yamato):

Substitute y = sqrt(x) then ^^

OpenStudy (anonymous):

ok will do

OpenStudy (unklerhaukus):

another way from here 8x/(1+x^1/2)=100 would be to rationalise the denominator ( 8x/(1+x^1/2) )×( (1-x^1/2)/(1-x^1/2) )=100

OpenStudy (anonymous):

I did rationalize the denominator and that's the point at which i was stuck. the result after rationalizing it was 8x -8x (x^1/2)/ (1-x) = 100

OpenStudy (anonymous):

This is the point I got to 8x-8x(x^1/2) = 100 - 100x 108x -8x (x^1/2) = 100 . How do i deal with the ever present (x^1/2) ?

OpenStudy (kira_yamato):

You could just substitute y as x^1/2 i.e. x = y^2

OpenStudy (anonymous):

\[108x -8x(\sqrt{x}) = 100\]

OpenStudy (kira_yamato):

\[108y^2 - 8y^3 = 100\]

OpenStudy (anonymous):

Currently working using that now.

OpenStudy (unklerhaukus):

i dont think there are nice solutions

OpenStudy (anonymous):

I don't think so either!

OpenStudy (anonymous):

108y^2 -8x^3 =100. This is difficult to solve!!!

OpenStudy (anonymous):

mistake 8y^3 not 8x^3

OpenStudy (unklerhaukus):

your on the right track,

OpenStudy (anonymous):

Thanks to all who tried to help! I will have to get back to it tomorrow!